This is a Template of How to Put Two ps-Made Pictures into a L^{A}TEX-Made Text and How to

Make the References

Innokentiy Semushin

Ulyanovsk State University, 42 Leo Tolstoy Str., 432970 Ulyanovsk, Russia

     SemoushinIV@ulsu.ru    http://staff.ulsu.ru/semoushin/

Attention! The proof that minimum of Auxiliary Performance Index is necessary and sufficient condition for filter optimality has been published in Englishlanguage literatute

in [8]. However in Russian, it was published earlier in [6] and [7].

A = [ \begin{matrix}0 & 1 \\ a_{1} & a_{2} \end{matrix} ] ,  D = [ \begin{matrix}d_{1} \\ d_{2} \end{matrix} ] ,  θ̂^{T} = [ a_{1} a_{2} d_{1} d_{2} ] .  (1)

References

1. Basseville, M., Nikiforov, I.: Detection of Abrupt Changes: Theory and Applications. Prentice-Hall Inc., New York (1993)

2. Landau, I.D. (ed.): Identification des Systemes. Les Bases de l’Identification des Systemes. Hermes, Paris (2001)

3. Caines, P.: Linear Stochastic Systems. John Willey & Sons, New York Chichester Brisbane Toronto Singapore (1988)

4. Soderstrom, T.: Identification of stochastic linear systems in presence of input noise. Automatica 17 (1981) 713–725

Fig. 1. Tracking the SSKF parameters k_{1} and k_{2} by d_{1} (left) and d_{2} (right) of (1)

5. Ansay, P., Gevers, M., Wertz, V.: Closed-loop or open-loop models in identification for control? In: Proc. of the European Control Conference (1999) CD-ROM F544

6. Semoushin, I.V.: Adaptive Identification and Fault Detection Methods in Random Signal Processing. Saratov University Publishers, Saratov (1985) [in Russian]

7. Semoushin, I.V.: Adaptive control for stochastic linear plants under conditions of uncertainty. Nonlinear Dynamical Systems - Qualitative Analysis and Control: Collected Papers, The Institute for System Analysis of the Russian Academy of Sciences 2 (1994) 104–110

8. Semoushin, I.V., Tsyganova, J.V.: Indirect error control for adaptive filtering. In: Neittaanmaki, P., Tiihonen, T., Tarvainen, P. (eds.): Proc. of the 3rd European Conference on Numerical Mathematics and Advanced Applications. World Scientific, Singapore New Jersey London Hong Kong (2000) 333–340

9. Semoushin, I., Gorokhov, O.: Computational processes in iterative control design. In: Sloot, P.M.A., Kenneth Tan, C.J., Dongarra, J.J., Hoekstra, A.G. (eds.): Computational Science – ICCS 2002. Lecture Notes in Computer Science, Vol. 2329. Springer-Verlag, Berlin Heidelberg New York Barcelona Hong Kong London Milan Paris Tokyo (2002) 186–195

10. Kondratiev, A.E., Fatyanova, O.A.: Start-stop algorithms for adaptive filtering. In: Boulyarski, S.V. (ed.): Young Researchers’ Collected Papers of the Ulyanovsk State University. UlSU Publishers, Ulyanovsk (2001) 6–7